# Projective rigidity of point-line configurations in the plane

### v1

@@ 3. Rigidity and flexibility of projective configurations
(3) $|I|=2|L|+2|P|-8\text{ and}$
(4) $|I^{\prime}|\leq 2|L(I^{\prime})|+2|P(I^{\prime})|-8,$
where the inequality must hold for all non-empty subsets of incidences $I^{\prime}\subseteq I$. An independent configuration, i.e. a configuration such that the rows of $M(S,\mathbf{p},\mathbf{l})$ are linearly independent has to satisfy (4).
$a$$b$$c$$l_{1}$$l_{2}$$p_{0}$$p_{1}$$p_{2}$$p_{3}$$p_{4}$$p_{5}$
Figure 1. The Pappus configuration – a $9_{3}$-configuration.
Example 3.2 (The Pappus configuration).
The Pappus configuration (see Figure 1) has $9$ points and $9$ lines, and as all lines are incident to three points, it has $3\times 9=27$ incidences. The Pappus configuration satisfies the count in (4). Furthermore, $|I|=27=2|P|+2|L|-9=2\times 9+2\times 9-9$, so one would predict that the Pappus configuration has a one-dimensional space of motions, after excluding the 8-dimensional space of projective transformations.
Suppose that the points $p_{0}$, $p_{2}$, $p_{3}$ and $p_{5}$ are pinned to exclude the projective transformations. A classic theorem of Pappus guarantees that $a$, $b$ and $c$ are collinear if the points $p_{1}$ and $p_{4}$ are on the lines $l_{1}=p_{0}p_{2}$ and $l_{2}=p_{3}p_{5}$, respectively. Hence $p_{1}$ and $p_{4}$ can be moved, independently, along $l_{1}$ and $l_{2}$, and $a$, $b$ and $c$ will remain collinear. The incidence that is implied by Pappus’ Theorem gives the configuration a two-dimensional space of motions, rather than the one-dimensional space of motions one might expect.
Example 3.3 (The Desargues configuration).
@@
The Desargues configuration has $10$ points, $10$ lines and $30$ incidences. So, $|I|=2|P|+2|L|-10$, which means the Desargues configuration is expected to have a two-dimensional space of motions, excluding the projective transformations.
@@
Figure 2. The $10_{3}$ Desargues configuration.
@@
The Pappus configuration and the Desargues configuration are both examples where counting the points, lines and incidences does not predict the correct number of motions, which means that the natural sparsity counts of Equations (3) and (4) given by counting the number of equations and variables in the rigidity matrix are not sufficient for characterizing minimal rigidity and independence in the rigidity matrix.
@@
Note that in both of these examples, it is a theorem in projective geometry that causes the discrepancy between the sparsity counts and the rank of the rigidity matrix. In Example 3.2, the incidence between $c$ and the line $ab$ is implied by Pappus’s theorem, and this causes the Pappus configuration to be more flexible than expected. Similarly, in the Desargues configuration, one of the incidences is implied by Desargues’s theorem, which causes the Desargues configuration to be more flexible than expected.
@@
We conjecture that this is part of a more general phenomenon, namely that for projective theorems regarding point-line configurations, where certain incidences imply others, these statements can be formulated in terms of the projective rigidity matrix having a non-trivial row dependence. Row dependencies, or self-stresses, are the subject of Section 5 in this paper.
4. Rigidity versus infinitesimal rigidity
4.1. Infinitesimal rigidity implies rigidity
Suppose that $(S,\mathbf{p},\mathbf{l})$ is a configuration of points and lines realizing an incidence geometry $S=(P,L,I)$. For simplicity move the configuration so that there are no points at infinity or lines through the origin.
In the following we will assume that four of the points (in general position) are pinned, which removes the trivial infinitesimal motions. Suppose also that we have a motion $\mathbf{p}_{i}(t)$ of each point $p_{i}$ and a motion $\mathbf{l}_{j}(t)$ of each line with $\mathbf{p}_{i}(t)\cdot\mathbf{l}_{j}(t)=0$ for all $t$ if $(p_{i},l_{j})\in I$, and $\mathbf{p}_{i}(t)=\mathbf{p}_{i}(0)$ for all $t$ if point $i$ is pinned.
If the configuration is rigid, then the only motion is zero not only at the pins but at all points. In general, if there is a motion, then the $n$-th derivative of the motion is $\sum_{k=0}^{n}{n\choose k}\mathbf{p}^{(k)}_{i}(t)\cdot\mathbf{l}^{(n-k)}_{j}(t)=0$. Evaluating at 00 gives $\sum_{k=0}^{n}{n\choose k}\mathbf{p}^{(k)}_{i}(0)\cdot\mathbf{l}^{(n-k)}_{j}(0)=0$, and for $k=1$, this says that
(5) $(\mathbf{p}^{(1)}_{1}(0),\ldots,\mathbf{l}^{(1)}_{1}(0),\ldots)\quad\forall\ (i,j)\in I$
is in the kernel of the rigidity matrix. A pinned configuration is infinitesimally rigid if the only element of the kernel of the pinned rigidity matrix, where the eight rows corresponding to $x_{i}^{(1)}=y_{i}^{(1)}=0$ for the four pinned points are added, is the zero vector.
Lemma 4.1.
Suppose that a pinned projective configuration is infinitesimally rigid. Then for all $n>0$, $\mathbf{p}_{i}^{(n)}(0)=\mathbf{0}$ for all points and $\mathbf{l}_{j}^{(n)}(0)=\mathbf{0}$ for all lines.
Proof.
The proof is by induction on $n$. Infinitesimal rigidity gives the base case, so let $n>1$ and suppose $\mathbf{p}_{i}^{(m)}(0)=\mathbf{0}$ for all points and $\mathbf{l}_{j}^{(m)}(0)=\mathbf{0}$ for all lines for all $m<n$. The sum $\sum_{k=0}^{n}{n\choose k}\mathbf{p}^{(k)}_{i}(0)\cdot\mathbf{l}^{(n-k)}_{j}(0)=0$ has only two non-zero terms by the inductive hypothesis, reducing to the system $\mathbf{p}_{i}(0)\cdot\mathbf{l}^{(n)}_{j}(0)+\mathbf{p}^{(n)}_{i}(0)\cdot\mathbf{l}_{j}(0)=0$ for for all incident pairs. So the $n$-th derivatives are a solution to the pinned infinitesimal rigidity system for the framework, and so the $n$-th derivatives must all be zero. ∎
It is an easy corollary of Lemma 4.1 that if a configuration is infinitesimally rigid, then it can have only constant analytic motions, and must therefore be rigid, by Theorem 3.1. We have therefore shown the following theorem.
Theorem 4.2.
Every infinitesimally rigid projective configuration is rigid.
So if the pinned projective rigidity matrix has zero kernel, then the projective configuration must be rigid. Equivalently, if the ordinary projective rigidity matrix has a kernel of dimension exactly $8$, then the configuration is rigid.
The following example shows the construction of an infinitesimally rigid configuration.
Figure 3. A cyclic $10_{3}$-configuration.
Example 4.3.
@@ 4.1. Infinitesimal rigidity implies rigidity
A combinatorial $v_{k}$-configuration is cyclic if there is a cyclic permutation of the points that sends the configuration to itself. The configuration shown in Figure 3 is a cyclic $10_{3}$ configuration with $10$ points and $10$ lines (non-isomorphic to the Desargues configuration). It can be drawn with 5-fold rotational symmetry and two symmetry classes of points and lines. Geometric configurations with this property are sometimes called astral configurations [12, Chapter 3]. This particular configuration can be constructed as follows [5]:
  1. (a)
@@
The $20_{4}$ configuration constructed this way is projectively infinitesimally rigid (as can be checked by computing the rank of the projective rigidity matrix), and therefore rigid, by Theorem 4.2.
Figure 4. A cyclic $20_{4}$-configuration.
4.2. Rigidity does not imply infinitesimal rigidity
Infinitesimal rigidity is a convenient linear certificate of rigidity, however, it is a strictly stronger requirement. To construct an infinitesimally rigid, but flexible configuration, we start with some inductive constructions. If $\mathbf{p}$ and $\mathbf{q}$ are the coordinates of two points (lines), then $[\mathbf{p},\mathbf{q}]$ denotes the line containing the two points (the point of intersection of the two lines).
Theorem 4.4.
Let $S=(P,L,I)$ be an independent configuration and suppose that $p_{i}$ and $p_{i^{\prime}}$ are distinct points with $[\mathbf{p}_{i},\mathbf{p}_{i^{\prime}}]\neq\mathbf{l}_{j}$ for any line $l_{j}$. Then adding a new line $l_{j}$ realized by $[\mathbf{p}_{i},\mathbf{p}_{i^{\prime}}]=\mathbf{l}_{j}$ and adding incidences $(p_{i},l_{j})$ and $(p_{i^{\prime}},l_{j})$ yields an independent configuration.
Dually, suppose that $l_{j}$ and $l_{j^{\prime}}$ are distinct lines with $[\mathbf{l}_{j},\mathbf{l}_{j^{\prime}}]\neq\mathbf{p}_{i}$ for any point $p_{i}$. Then adding a new point $p_{i}$ realized by $[\mathbf{l}_{j},\mathbf{l}_{j^{\prime}}]=\mathbf{p}_{i}$ and adding incidences $(p_{i},l_{j})$ and $(p_{i},l_{j^{\prime}})$ yields an independent configuration.
If the original configuration was infinitesimally rigid, then so is the new one.
Proof.
We need to show independence. By the homogeneity of the projective plane, we may assume that the points are all finite, and that the line through any pair of points does not pass through the origin, so that the matrix is defined. We may also assume that $\mathbf{p}_{i}$ and $\mathbf{p}_{i^{\prime}}$ are linearly independent. The only non-zero entries of the two columns associated to the line $l_{j}$ are in the rows for the incidences $(p_{i},l_{j})$ and $(p_{i^{\prime}},l_{j})$, and since we assumed that $\mathbf{p}_{i}$ and $\mathbf{p}_{i^{\prime}}$ are linearly independent, any row dependence must be zero on those rows. By assumption there is no row dependence of the remaining rows, so the configuration must be independent.
The second claim follows by duality, and infinitesimal rigidity follows from the rank of the new matrix, which is $|I|+2$, and $|I|+2=2|P|+2(|L|+1)-8$. ∎
A dyadic rational is a rational number whose denominator is a power of $2$. In the next theorem, we construct isostatic grids with points whose coordinates are dyadic rationals.
Theorem 4.5.
For each $N\geq 0$ there exists an independent, infinitesimally rigid configuration $S=(P,L,I)$ of points and lines such that every geometric point/line incidence between $\mathbf{p}_{i}$ and $\mathbf{l}_{j}$ in the plane is recorded as an incidence $(i,j)\in I$, and such that all points with coordinates $(n/2^{N}:m/2^{N}:1)$, for integers $n$, $m$ satisfying $0\leq n,m\leq 2^{N}$, is either a point of the configuration or is the point of intersection of at least two lines.
Proof.
We construct an example for each $N$ recursively.
At the 00-th level, first take the four points of the unit square $a=(0:0:1)$, $b=(0:1:1)$, $c=(1:1:1)$, and $d=(1:0:1)$. Then form the complete quadrilateral by constructing the lines $ab$, $ac$, $bc$ and $cd$ and the intersection $p_{1}=(1:0:0)$ of the lines $ad$ and $bc$, and the intersection $p_{2}=(0:1:0)$ of the lines $ab$ and $cd$. Note that the complete quadrangle is independent, and infinitesimally rigid.
Then add the diagonal lines, $ac$ and $bd$, and their intersections $p_{3}=(1:1:0)$ and $p_{4}=(1:-1:0)$ with the line through the points $p_{1}$ and $p_{2}$, which is the line at infinity. Then construct the lines $ap_{4}$, $bp_{3}$, $cp_{4}$ and $dp_{3}$. Then construct the points $e$ and $f$ as the intersections of the lines $ap_{4}$ and $bp_{3}$ and the lines $bp_{3}$ and $cp_{4}$ respectively.
Finally, construct the horizontal line through $e$, which is the line $ep_{1}$, and the vertical line through $f$, which is the line $fp_{2}$. See Figure 5a.
a) Refer to caption  b) Refer to caption
Figure 5. Isostatic dyadic grids for $N=0$ and $N=1$.
This completes the 00-th level, and yields an independent and infinitesimally rigid configuration by Theorem 4.4.
Notice for the purpose of the recursion that the segments $ab$ and $cd$ have points with dyadic coordinates at level $N$ and crossings, by single lines, exactly at heights $(2k+1)/2^{N+1}$, with $10\leq k\leq 2^{N}-$ 1. Notice also that segments $ae$, $eb$, $bf$ and $fc$ have points exactly those with dyadic coordinates at level $N+1$, and are only crossed by lines at those points.
For the recursive step start by the points of intersection of $ab$ and $cd$ with the horizontal lines of the configuration. Add the lines between each of the new points and $p_{3}$ and $p_{4}$. Each such new line bisects one of the segments along $ae$, $eb$, $bf$ and $fc$. Add points at each of these intersections. Finally, add horizontal lines through the new points along $ae$ and $eb$ and vertical lines through the new points along $bf$ and $fc$. These subdivide the grid, with the points along $ab$ and $cd$ again alternating between points and pairs of crossed lines.
The dyadic grids at levels $N=0$ and $N=1$ are illustrated in Figures 5b and 5a. Each level is created from the previous by intersections of two lines and drawing lines through two points, and so each level is independent and infinitesimally rigid by Theorem 4.4. ∎
In the construction, the points at level $N$, i.e. the points with coordinates $(n/2^{N}:m/2^{N}:1)$, in the interior of $abcd$ occur at the crossings of four lines. If a point is introduced there by the Cayley algebra, then that point will have two incidences recorded in the set $I$, and two implied geometric incidences. Adding either of these geometric incidences to $I$ results in a loss of constraint independence, in other words, the introduction of an equilibrium stress. This illustrates the connection between self-stresses, constraint independence and projective theorems: certain incidences are implied by projective theorems, and such implied incidences will be dependent of the remaining incidences.
The points at the level $N+1$ inside $abcd$ which occur at the crossings of two lines are also the points of a square lattice, and all of these points may be added to form a rigid and independent configuration. See Figure 5b.
Corollary 4.6.
There exists an independent, infinitesimally rigid configuration $S=(P,L,I)$ in which every geometric point/line incidence between $\mathbf{p}_{i}$ and $\mathbf{l}_{j}$ corresponds to an incidence in $(i,j)\in I$, and containing as many points of a square lattice as required.
a) Refer to caption  b) Refer to caption
Figure 6. An isostatic configuration with $40$ lattice points.
The isostatic configuration constructed in Corollary 4.6 can be further used to construct the example in the following theorem, which shows that infinitesimal rigidity does not imply rigidity.
Theorem 4.7.
There exists a configuration which is rigid but not infinitesimally rigid.
Proof.
This proof uses the classical method of constructing points on a quadratic curve given five points.
Let there be a projective configuration as in the following figure with the white vertices pinned, from which the black points and lines can be constructed and added to the rigid configuration. The white points can be constructed, and pinned, as part of an isostatic grid, as in Corollary 4.6. This can be done starting from four arbitrary points. See Figure 7.
Figure 7. A projective mechanism, pinned at the white points, whose motion requires point $f$ to move along the red hyperbola.
Now, add a new point $t$, in green, incident to $oy$, creating a one-degree of freedom mechanism. The point $t$ moves freely on the fixed line $oy$, and the configuration will remain a mechanism if we add line $jt$, which we will think of as the Pascal line of the hexagon $abcdef$ which appears after constructing the additional green points and lines; the intersection $i$ of the lines $cd$ and $jt$, the intersection $k$ of the lines $bc$ and $jt$, the lines $ai$ and $ek$, and finally the intersection $f$ of the lines $ai$ and $ek$.
Each step preserves the mechanism, and forces the point $f$ to move on the hyperbola determined by the pinned points $\{a,b,c,d,e\}$. Finally, require an incidence between $f$ and the line $ox$. Consider what happens if we engineer this configuration so that the line $ox$ cuts the hyperbola in exactly one point, a point of tangency to the hyperbola through the points $\{a,b,c,d,e\}$. Care must be taken to get the original pinning so that the tangency can be achieved.
Now the resulting configuration is rigid, and there is a unique realization which satisfies the final required incidence. We will show that this configuration is not infinitesimally rigid. A velocity at $t$ along $oy$ preserves the mechanism, and has consequential velocities which are zero on all but the yellow points and lines. These velocities must be in the kernel of the incidence rigidity matrix, so, we know that the consequential velocity at $f$ must lie along the pinned line $ox$. See Figure 8.
Figure 8. A rigid pinned configuration with an infinitesimal motion.
in which the velocities at $i$ and $k$, easily seen to be non-zero if $v_{t}$ is non-zero, are not shown. All green points move infinitesimally or none. We know that there is a finite motion without the final incidence which is consistent with the constraint of the last incidence, so this velocity assignment is a non-trivial infinitesimal motion of the configuration. ∎
This theorem shows that one cannot in general establish the non-rigidity of a configuration by considering the constraint matrix alone. Furthermore, in the traditional setting of bar-and-joint frameworks, infinitesimal rigidity and rigidity are equivalent for generic frameworks. This means that rigid but not infinitesimally rigid bar-and-joint framework have to be in some special position. The configuration in Theorem 4.7, however, can be constructed starting from four arbitrary points, so it is not in a special geometric position.
5. Self-stresses of projective configurations
The theory of infinitesimal rigidity (column dependencies) is a dual theory of row dependencies which engineers identify as “statics”. Statics generally gives a different set of tools and insights.
@@ 5. Self-stresses of projective configurations
(6) $\sum_{l_{j}\in L}\omega_{i,j}\mathbf{l}_{j}=\mathbf{0},\qquad\sum_{p_{i}\in P}\omega_{i,j}\mathbf{p}_{i}=\mathbf{0}.$
@@
Define new stress coefficients by $\omega_{i,j}^{\prime}=\lambda_{i}\lambda_{j}\omega_{i,j}$ for each incidence $(p_{i},l_{j})$. Now consider the sum $\Sigma_{l_{j}\in L}\omega^{\prime}_{i,j}\mathbf{l}^{\prime}_{j}$. By definition $\Sigma_{l_{j}\in L}\omega^{\prime}_{i,j}\mathbf{l}^{\prime}_{j}=\Sigma_{l_{j}\in L}\lambda_{i}\lambda_{j}\omega_{i,j}\mathbf{l}^{\prime}_{j}$, and by assumption $\Sigma_{l_{j}\in L}\lambda_{i}\lambda_{j}\omega_{i,j}\mathbf{l}^{\prime}_{j}=\lambda_{i}A(\Sigma_{l_{j}\in L}\omega_{i,j}\mathbf{l}_{j})=\mathbf{0}$. Similarly, $\Sigma_{p_{i}\in P}\omega^{\prime}_{i,j}\mathbf{p}^{\prime}_{i}=\lambda_{j}A(\Sigma_{p_{i}\in P}\omega_{i,j}\mathbf{p}_{i})=\mathbf{0}$.
Hence $\omega^{\prime}_{i,j}$ are an equilibrium stress on the realization $(S,\mathbf{p^{\prime}},\mathbf{l^{\prime}})$. ∎
@@
(7) $\sum_{l_{j}\in L}\omega_{i,j}(a_{j},b_{j})=\mathbf{0},\qquad\sum_{p_{i}\in P}\omega_{i,j}(x_{i},y_{i})=\mathbf{0}$
hold for all points $p_{i}\in P$ and lines $l_{j}\in L$. Equation (6) is satisfied for a point $p_{i}$ if Equation (7) is satisfied, and the sum $\Sigma_{l_{j}:(p_{i},l_{j})\in I}w_{i,j}$, which gives the third coordinate of Equation (6), is zero. Dually, for lines, Equation (7) has to hold, and the sum $\Sigma_{p_{i}:(p_{i},l_{j})\in I}w_{i,j}$ has to be 0. The equations $\sum_{l_{j}\in L:(p_{i},l_{j})\in I}\omega_{i,j}=0$ and $\sum_{p_{i}\in P:(p_{i},l_{j})\in I}\omega_{i,j}=0$ say that, for a configuration placed with no points at infinity or lines through the origin, every stress $\omega$ must be purely combinatorial, that is, it must be a row dependence of the combinatorial incidence matrix.
@@
The dual statement holds for lines. Hence, a row-dependence in the rigidity matrix implies that the configuration has an equilibrium stress.
@@
The conditions that the row space of $M(S,\mathbf{p},\mathbf{l})$ has rank $82|P|+2|L|-$ 8, i.e. the configuration is infinitesimally rigid, is equivalent to the condition of the configuration being static rigidity, which says that the vectors of the from $\omega M(S,\mathbf{p},\mathbf{l})$ span the entire orthogonal complement of the space of trivial motions.
@@
Figure 9. The complete quadrilateral.
Example 5.2 (The complete quadrilateral).
@@
Every point of the complete quadrilateral is incident to exactly two lines. If the point $p_{i}$ lies on the lines $l_{j}$ and $l_{k}$, then $\omega_{i,j}\mathbf{l}_{j}+\omega_{i,k}\mathbf{l}_{k}=0$. As we are interested in the case where $\omega_{j,i}$ and $\omega_{k,i}$ are non-zero, we must have that $\mathbf{l}_{j}$ is a scalar multiple of $\mathbf{l}_{k}$. Geometrically, this means that $\mathbf{l}_{j}$ and $\mathbf{l}_{k}$ are the same line. Because the complete quadrilateral is connected, this implies that all lines have to be the same in a self-stressed realization.
@@
(8) $\sum_{j:(i,j)\in E}s_{ij}(x_{ij},y_{ij},1)=\mathbf{0}.$
In Example 5.2, we find that a dependence in the projective matrix includes a check that the scalars, restricted to the columns for the lines, form a self-stress for the weaving of the lines. See [20, 19]. The converse does not hold.
6. Symmetric projective configurations
Many examples of projective configurations exhibit symmetry. There are several reasons for this; symmetry implies beauty, and usually the symmetry makes it easier to construct the configuration. For example, the $21_{4}$ Grünbaum-Rigby configuration, whose description in 1990 [13] began the modern study of configurations, is typically drawn with 7-fold dihedral symmetry. Many papers have constructed interesting examples of configurations (for example, [7, 5, 2, 3, 4, 1], and [12] has many other examples), including those shown in Figures 3 and 4, by leveraging symmetry and geometry to prove that the necessary incidences occur.

### v3

@@ 3 Rigidity and flexibility of projective configurations
$|I|=2|L|+2|P|-8\text{ and}$ (4)
$|I^{\prime}|\leq 2|L(I^{\prime})|+2|P(I^{\prime})|-8,$ (5)
where the inequality must hold for all non-empty subsets of incidences $I^{\prime}\subseteq I$. A configuration such that the rows of $M(S,\mathbf{p},\mathbf{l})$ are linearly independent has to satisfy (5). Such configurations are said to be independent.
Example 1 (The Desargues configuration).
@@
The Desargues configuration has $10$ points, $10$ lines and $30$ incidences. So, $|I|=2|P|+2|L|-10$, which means that the Desargues configuration is expected to have a two-dimensional space of motions, excluding the projective transformations.
@@
Figure 1: The $10_{3}$ Desargues configuration.
@@
The Desargues configuration is an example where counting the points, lines and incidences does not predict the correct number of motions, which means that the natural sparsity counts of Equations (4) and (5) are not sufficient for characterizing independence in the rigidity matrix and minimal rigidity.
@@
Note that it is a theorem in projective geometry that causes the discrepancy between the sparsity counts and the rank of the rigidity matrix. In the Desargues configuration, one of the incidences is implied by Desargues’s theorem, which causes the Desargues configuration to be more flexible than expected.
@@
We conjecture that this is part of a more general phenomenon, namely that for projective theorems regarding point-line configurations, where certain incidences imply others, these statements can be formulated in terms of the projective rigidity matrix having a non-trivial row dependence. More evidence for this conjecture can be found in [6]. Row dependencies, or self-stresses, are the subject of Section 4 in this paper.
We conclude this section with an example of an infinitesimally rigid configuration.
Figure 2: A cyclic $10_{3}$-configuration.
Example 2.
@@
A combinatorial $v_{k}$-configuration is cyclic if there is a cyclic permutation of the points that sends the configuration to itself. The configuration shown in Figure 2 is a cyclic $10_{3}$ configuration with $10$ points and $10$ lines (non-isomorphic to the Desargues configuration). It can be drawn with 5-fold rotational symmetry and two symmetry classes of points and lines. Geometric configurations with this property are sometimes called astral configurations [15, Chapter 3]. This particular configuration can be constructed as follows [5]:
  1. 1.
@@
The $20_{4}$ configuration constructed this way is projectively infinitesimally rigid (as can be checked by computing the rank of the projective rigidity matrix), and therefore rigid, by Theorem 2.
Figure 3: A cyclic $20_{4}$-configuration.
4 Self-stresses of projective configurations
In this section, we are going to introduce the theory of row dependencies, or self-stresses. The theory of row dependencies is a dual theory of column dependencies (infinitesimal motions).
@@ 4 Self-stresses of projective configurations
$\sum_{l_{j}\in L}\omega_{i,j}\mathbf{l}_{j}=\mathbf{0},\qquad\sum_{p_{i}\in P}\omega_{i,j}\mathbf{p}_{i}=\mathbf{0}.$ (6)
@@
Define new stress coefficients by $\omega_{i,j}^{\prime}=\lambda_{i}\lambda_{j}\omega_{i,j}$ for each incidence $(p_{i},l_{j})$. Now consider the sum $\Sigma_{l_{j}\in L}\omega^{\prime}_{i,j}\mathbf{l}^{\prime}_{j}$. By definition,
$\Sigma_{l_{j}\in L}\omega^{\prime}_{i,j}\mathbf{l}^{\prime}_{j}=\Sigma_{l_{j}\in L}\lambda_{i}\lambda_{j}\omega_{i,j}\mathbf{l}^{\prime}_{j}$
and by assumption,
$\Sigma_{l_{j}\in L}\lambda_{i}\lambda_{j}\omega_{i,j}\mathbf{l}^{\prime}_{j}=\lambda_{i}A(\Sigma_{l_{j}\in L}\omega_{i,j}\mathbf{l}_{j})=\mathbf{0}.$
Similarly,
$\Sigma_{p_{i}\in P}\omega^{\prime}_{i,j}\mathbf{p}^{\prime}_{i}=\lambda_{j}A(\Sigma_{p_{i}\in P}\omega_{i,j}\mathbf{p}_{i})=\mathbf{0}.$
Hence $\omega^{\prime}_{i,j}$ is an equilibrium stress on the realization $(S,\mathbf{p^{\prime}},\mathbf{l^{\prime}})$. ∎
@@
$\sum_{l_{j}\in L}\omega_{i,j}(a_{j},b_{j})=\mathbf{0},\qquad\sum_{p_{i}\in P}\omega_{i,j}(x_{i},y_{i})=\mathbf{0}$ (7)
hold for all points $p_{i}\in P$ and lines $l_{j}\in L$. Equation (6) is satisfied for a point $p_{i}$ if Equation (7) is satisfied, and the sum $\Sigma_{l_{j}:(p_{i},l_{j})\in I}w_{i,j}$, which gives the third coordinate of Equation (6), is zero. Dually, for lines, Equation (7) has to hold, and the sum $\Sigma_{p_{i}:(p_{i},l_{j})\in I}w_{i,j}$ has to be 0. The equations $\sum_{l_{j}\in L:(p_{i},l_{j})\in I}\omega_{i,j}=0$ and $\sum_{p_{i}\in P:(p_{i},l_{j})\in I}\omega_{i,j}=0$ say that, for a configuration placed with no points at infinity or lines through the origin, every self-stress $\omega$ must be purely combinatorial, that is, it must be a row dependence of the combinatorial incidence matrix.
@@
The dual statement holds for lines. Hence, a row dependence in the rigidity matrix implies that the configuration has an equilibrium stress.
@@
The conditions that the row space of $M(S,\mathbf{p},\mathbf{l})$ has rank $82|P|+2|L|-$ 8, i.e. the configuration is infinitesimally rigid, is equivalent to the condition of the configuration being statically rigid, which says that the vectors of the from $\omega M(S,\mathbf{p},\mathbf{l})$ span the entire orthogonal complement of the space of trivial motions.
@@
Figure 4: The complete quadrilateral.
Example 3 (The complete quadrilateral).
@@
Every point of the complete quadrilateral is incident to exactly two lines. If the point $p_{i}$ lies on the lines $l_{j}$ and $l_{k}$, then $\omega_{i,j}\mathbf{l}_{j}+\omega_{i,k}\mathbf{l}_{k}=0$. As we are interested in the case where $\omega_{j,i}$ and $\omega_{i,k}$ are non-zero, we must have that $\mathbf{l}_{j}$ is a scalar multiple of $\mathbf{l}_{k}$. Geometrically, this means that $\mathbf{l}_{j}$ and $\mathbf{l}_{k}$ are the same line. Because the complete quadrilateral is connected, this implies that all lines have to be the same in a self-stressed realization.
@@
$\sum_{j:(i,j)\in E}s_{ij}(x_{ij},y_{ij},1)=\mathbf{0}.$ (8)
In Example 3, we find that a dependence in the projective matrix includes a check that the scalars, restricted to the columns for the lines, form a self-stress for the weaving of the lines. See [25, 24]. The converse does not hold.
5 Symmetric projective configurations
Many examples of projective configurations exhibit symmetry. There are several reasons for this; symmetry implies beauty, and usually the symmetry makes it easier to construct the configuration. For example, the $21_{4}$ Grünbaum-Rigby configuration, whose description in 1990 [16] began the modern study of configurations, is typically drawn with 7-fold dihedral symmetry. Many papers have constructed interesting examples of configurations (for example, [8, 5, 2, 3, 4, 1], and [15] has many other examples), including those shown in Figures 2 and 3, by leveraging symmetry and geometry to prove that the necessary incidences occur.